OpenStudy (anonymous):

what is the slope of a line that is parallel to the line represented by the equation y=-2/5x+4? ty

OpenStudy (anonymous):

a line is always represented by the following equation: Ax+By+C=0 its slope is giving by: m=-A/B You have to put you equation into that form too 2/5x+y-4=0 so the slope for this like is: m=-2/5/1 m=-2/5

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

np